OpenStudy (anonymous):

How would I solve 75+3x^2?

OpenStudy (anonymous):

perhaps you mean 75+3x^2 = 0?

hartnn (hartnn):

to solve u need some equality....

OpenStudy (anonymous):

yes thats what I mean

OpenStudy (anonymous):

Ok, rearrange to get x on its own: 3x^2 = -75 x^2 = -25 x = √-25 which is not a real number

OpenStudy (anonymous):

Did you mean 75x+3x^2 = 0?

OpenStudy (lgbasallote):

let's just mean everything shall we

OpenStudy (lgbasallote):

i think it's just asking to factor out 75 + 3x^2

OpenStudy (anonymous):

Start with the given expression Factor out the GCF 3 Now let's focus on the inner expression ------------------------------------------------------------ Looking at we can see that the first term is and the last term is where the coefficients are 1 and 25 respectively. Now multiply the first coefficient 1 and the last coefficient 25 to get 25. Now what two numbers multiply to 25 and add to the middle coefficient 10? Let's list all of the factors of 25: Factors of 25: 1,5 -1,-5 ...List the negative factors as well. This will allow us to find all possible combinations These factors pair up and multiply to 25 1*25 5*5 (-1)*(-25) (-5)*(-5) note: remember two negative numbers multiplied together make a positive number Now which of these pairs add to 10? Lets make a table of all of the pairs of factors we multiplied and see which two numbers add to 10

OpenStudy (lgbasallote):

lol looks the same answer as the one here http://www.algebra.com/algebra/homework/Polynomials-and-rational-expressions/Polynomials-and-rational-expressions.faq.question.142171.html i think @jim_thompson5910 wrote it =)))

OpenStudy (lgbasallote):

and fyi @justindrake that problem is so different from this one =))))

Parth (parthkohli):

What a coincidence! Wow! Jim_Thompson!

OpenStudy (anonymous):

@lgbasallote My Answer Is Right

OpenStudy (anonymous):

the problem reads 75+3x^2=0. so i move the 3x^2 to the other side is that what you said?

OpenStudy (lgbasallote):

*your* answer?

OpenStudy (anonymous):

@lgbasallote yeah

OpenStudy (anonymous):

What? ^

OpenStudy (anonymous):

So go with @justindrake's answer?

OpenStudy (anonymous):

yeah my ans is right

OpenStudy (anonymous):

Okay is it the one where you typed the paragraph?

OpenStudy (anonymous):

Can you type how I'm supposed to start it plz?

OpenStudy (farmdawgnation):

@justindrake Please don't copy and paste answers from other sites and make small tweaks. Thanks.

OpenStudy (anonymous):

@farmdawgnation im not copying and pasteing the ans from website im copying it from my brain i dont know that this site have ans to